OpenStudy (anonymous):

convert the number to polar form 4(cos60º + i sin60º)

terenzreignz (terenzreignz):

<scratches head> You mean rectangular form, right?

OpenStudy (anonymous):

\[\Large 4 e^{i\frac \pi 3} \]

OpenStudy (anonymous):

how do i get pi/3? :)

OpenStudy (anonymous):

180/3=60

OpenStudy (anonymous):

\[ \pi \, \text{radians} = 180^{\circ} \]

OpenStudy (anonymous):

ok :) do you take the angle in cos, or sin ?

OpenStudy (anonymous):

It has to be the same angle

OpenStudy (anonymous):

ok thanks !!

OpenStudy (anonymous):

YW